# The Expressive Power of Transformers with Chain of Thought

### 20 Mar 2024 v4

@@ Abstract
Recent theoretical work has identified surprisingly simple reasoning problems, such as checking if two nodes in a graph are connected or simulating finite-state machines, that are provably unsolvable by standard transformers that answer immediately after reading their input. However, in practice, transformers’ reasoning can be improved by allowing them to use a “chain of thought” or “scratchpad”, i.e., generate and condition on a sequence of intermediate tokens before answering. Motivated by this, we ask: Does such intermediate generation fundamentally extend the computational power of a decoder-only transformer? We show that the answer is yes, but the amount of increase depends crucially on the amount of intermediate generation. For instance, we find that transformer decoders with a logarithmic number of decoding steps (w.r.t. the input length) push the limits of standard transformers only slightly, while a linear number of decoding steps, assuming a slight generalization to standard pre-norm, adds a clear new ability (under standard complexity conjectures): recognizing all regular languages. Our results also imply that linear steps keep transformer decoders within context-sensitive languages, and polynomial steps with generalized pre-norm make them recognize exactly the class of polynomial-time solvable problems—the first exact characterization of a type of transformers in terms of standard complexity classes. Together, our results provide a nuanced framework for understanding how the length of a transformer’s chain of thought or scratchpad impacts its reasoning power.
@@ 1 Introduction
In this work, we address these open questions by characterizing the reasoning power of transformer decoders that can take intermediate steps before generating an answer and comparing them against transformers without intermediate steps. A transformer with a chain of thought constitutes a special case of a transformer decoder with intermediate steps. Our fine-grained results give upper and lower bounds on transformers’ power depending on $t(n)$: the number of allowed intermediate steps as a function of the input size $n$. We focus mainly on understanding three regimes: logarithmic steps (when $t(n)=\Theta(\log n)$), linear steps (when $t(n)=\Theta(n)$), and polynomial steps.
@@
Linear Steps. Linear intermediate steps allow transformers with generalized pre-norm to simulate automata ($\mathsf{NC}^{1}$-complete), which cannot be done without intermediate steps unless $\mathsf{TC}^{0}=\mathsf{NC}^{1}$.
Polynomial Steps. With a polynomial number of decoding steps, we show that transformers with strict causal attention and generalized pre-norm are equivalent to the class $\mathsf{P}$. This is, to our best knowledge, the first equivalence between a class of transformers and a standard complexity class.
Together, our results provide a framework for understanding how the length of a transformer’s chain of thought or scratchpad affects its reasoning power. We find that a logarithmic chain does not add much, while a linear chain affords more power on inherently sequential reasoning problems, compared to transformers with no chain of thought.
@@ 1.1 Main Results: Power of Transformers with Intermediate Decoding
Here $\mathsf{CoT}(t(n))$ denotes the set of languages recognized by some transformer using $t(n)$ decoding steps. Our lower bound (left side of Equation 1) assumes strict causal saturated attention and generalized pre-norm, while upper bounds hold both with and without these architectural assumptions. Both our time lower bound and space upper bound are fairly tight: improving either by a factor larger than $\log t(n)$ would result in a fundamental complexity theory advance (Hopcroft et al., 1977).
@@
The left side of Equation 1 is proven by showing transformer decoders can simulate $t$ Turing machine steps with $t$ intermediate steps. Similar prior results have assumed a transformer with external memory (Schuurmans, 2023) or an encoder-decoder model with nonstandard-positional decodings (Pérez et al., 2021). Our construction adapts these ideas to work for a decoder-only model without external memory or extra positional encodings, but with strict causal masking and generalized pre-norm (cf. Section 2.1). [^4] Our construction (Theorem 2) can be easily modified to work with an encoder-decoder model as well. The key idea behind our more general construction is the layer-norm hash (Section 3.1): a simple module for effectively storing memory in decoder-only transformers. We believe the layer-norm hash could be broadly useful for building algorithms in transformers. For example, Yao et al. (2021) used a related idea to construct transformers that recognize bounded-depth Dyck languages, although in a more ad hoc way.
@@ 2.1 Transformers
    For each $1\leq\ell\leq d$ and $1\leq k\leq h$, a head similarity function $s^{\ell}_{k}:\mathbb{D}_{p}^{m}\times\mathbb{D}_{p}^{m}\to\mathbb{D}_{p}$ whose form is defined in Section A.2 (and includes generalized layer-norm);
@@
    For each $1\leq\ell\leq d$ and $1\leq k\leq h$, a head value function $v^{\ell}_{k}:\mathbb{D}_{p}^{m}\to\mathbb{D}_{p}^{m/h}$ whose form is defined in Section A.2 (and includes generalized layer-norm);
@@
    For each $1\leq\ell\leq d$, an activation function $f^{\ell}:(\mathbb{D}_{p}^{m/h})^{h}\times\mathbb{D}_{p}^{m}\to\mathbb{D}_{p}^{m}$ whose form is defined in Section A.3 and implicitly uses the feedforward dimension $w$ (and includes generalized layer-norm);
@@
    Multi-Pre-Norm. Our lower bound constructions require $s_{\ell}$ and $f_{\ell}$ in Definition 2 to allow a generalization of standard pre-norm. In standard pre-norm, each sublayer reads from the residual stream by applying a single layer-norm. We generalize so each sublayer takes $k$ different linear projections from the residual stream, applies layer-norm to each, and concatenates (cf. Definition 4 in Appendix A). Since it resembles multihead self-attention, we call this multi-pre-norm.
@@ 2.2 Models of Computation
We define computation with a Turing machine on input string $\sigma\in\Sigma^{*}$ as follows. A configuration of a Turing machine is a finite state $q$ along with the contents of an input tape $c^{0}$, $k$ work tapes $c^{1},\ldots,c^{k}$, and an output tape $c^{k+1}$. Finally, for each tape $\tau$, a configuration specifies a head position $h^{\tau}$. We start with the initial state $q_{0}$ and the input tape $c^{0}_{0}$ containing $\sigma$ starting at position 00 with infinite $b$’s on each side, and $h^{0}_{0}=0$. All other tapes start containing all $b$’s and with their head at 00. At each time step $i$, if $q_{i}\not\in F$, we recurrently update the configuration by first computing:
@@ 3 Lower Bounds for Transformer Decoders
Prior work (Merrill & Sabharwal, 2023a) has established strong upper bounds on the reasoning problems transformers can solve. Specifically, under standard conjectures in complexity, transformers without intermediate decoding cannot recognize all regular languages. In this section, we show some of these shortcomings can be overcome with a suitable number of intermediate decoding steps (and multi-pre-norm). Specifically, a linear number of steps enables simulating an automaton. We also show this can be extended to simulate $t(n)$ Turing machine steps with $t(n)$ decoding steps.
@@ 3.1 Introducing Layer-Norm Hash
$\phi(x,y)$ is a unit vector in $\mathbb{R}^{4}$. A key feature is scale invariance, and, in particular, that $\phi(x/i,1/i)$ is invariant w.r.t. $i$ in the sense that it is only a function of $x$, independent of $i$. Let $\phi_{x}\triangleq\phi(x,1)$. Then we have the following properties, whose proof may be found in Appendix B.
@@ 3.2 Simulating Automata
For any regular language $L$. there is a decoder-only multi-pre-norm transformer with strict causal saturated attention (with or without positional encodings) that, on input \$$\$x$, [^6] Theorem 1 goes through without strict causal masking (but Theorem 2 will require strict masking). [^7] Theorems 1 and 2 both go through without \$currency-dollar\\$\$ as long as token $j$ can compute value $v_{j}=\mathbbm{1}[j=0]$. checks whether $x\in L$ with $\left\lvert x\right\rvert+1$ decoding steps.
@@
Let $A$ be a finite state automaton recognizing $L$. The idea is to simulate one step of $A$ with one transformer decoding step (after first reading $n$ input tokens). We refer to tokens with 00 indexing: \$currency-dollar\\$\$ is token 0, $x_{1}$ is token 1, etc. At step $i,n\leq i\leq 2n$, we will output a token $q_{i-n}$ encoding the next state of $A$. After printing the final state $q_{n}$, we use one additional step to output $1$ iff $q_{n}\in F$, the set of final states of $A$. At each token $i>0$, we compute $1/i$ by attending uniformly over the strict left context with value $\mathbbm{1}[x_{j}=\$]$. We show by induction that at step $i\geq n$, we can output $q_{i-n}$.
@@
Inductive Case: $i>n$. We already have a sequence of intermediate tokens $q_{0},\ldots,q_{i-n-1}$. Our goal is to compute $q_{i-n}=\delta(q_{i-n-1},\sigma_{i-n})$, which first involves retrieving $q_{i-n-1}$ and $\sigma_{i-n}$. $q_{i-n-1}$ is the input to the current column of the transformer. We will use hard attention to retrieve the current input symbol $\sigma_{i-n}$. To do this, we attend uniformly over the prior decoding tokens and \$, with a value of $1$ at \$ and 00 elsewhere. At tokens $i>n$ (i.e., decoding tokens), this yields $\frac{1}{i-n}$. We then leverage the multi-pre-norm architecture to pass two layer-norms to a feedforward network:
@@
Theorem 1 shows that a linear number of decoding steps gives additional reasoning power to log-precision transformers with multi-pre-norm (assuming $\mathsf{TC}^{0}\neq\mathsf{NC}^{1}$). This follows because log-precision transformers with no decoding steps are contained in uniform $\mathsf{TC}^{0}$ (Merrill & Sabharwal, 2023b), which means they cannot recognize all regular languages. In contrast, Theorem 1 says a linear number of steps is sufficient for recognizing all regular languages, establishing a conditional separation. This is an example of simple and familiar additional computational power granted by additional decoding steps. The core challenge in simulating an automaton is recurrence, which cannot be done without decoding steps (Merrill & Sabharwal, 2023b). A linear number of decoding steps allows simulating recurrence, which is where the additional power comes from. However, this added power does not stop with finite-state machines: the layer-norm hash can be used to simulate more complex models of computation such as Turing machines, which we will turn to next.
@@ 3.3 Simulating Turing Machines
We now show how a transformer decoder can simulate a Turing machine in real time using the layer-norm hash. Our decoder-only construction resembles the encoder-decoder construction of Pérez et al. (2021). However, it avoids simplifying assumptions from Pérez et al. (2021). In addition to assuming non-standard attention and no layer-norm, they required $1/i,1/i^{2},$ and $i$ in the positional embeddings, which is problematic because transformers cannot represent unbounded scalars like $i$ due to layer-norm. In contrast, our construction works with or without positional encodings. However, it assumes strict causal masking and multi-pre-norm (Section 2.1).
@@
Let $M$ be a Turing machine that, on input length $1+n$, runs for at most $t(n)$ steps (at most polynomial). There is a decoder-only multi-pre-norm transformer with strict causal saturated attention (with or without positional encodings) that, on input \$$\$x$, takes $t(n)$ decoding steps and then, with $\left\lvert M(x)\right\rvert$ additional steps, outputs $M(x)$.
@@
Inductive Case: $i>n$. We first use a layer to reconstruct each $h^{\tau}_{i}$, the current position on tape $\tau$ at step $i$. For each $\tau$, a head attends with query $1$, key $\mathbbm{1}[x_{j}\not\in\Sigma]$, and value being the move direction of $\tau$ at $j$. Since we assume strict causal attention (for reasons that will become clear later), head $\tau$ thus computes $h^{\tau}_{i-1}/i$. Since we need $h^{\tau}_{i}$, we write both $(h^{\tau}_{i-1}\pm 1)/i$ to the residual stream; recall that we had computed $1/i$ earlier. Whenever we need $h^{\tau}_{i}/i$ going forward, we use a linear layer to select either $(h^{\tau}_{i-1}+1)/i$ or $(h^{\tau}_{i-1}-1)/i$ depending on whether the current input $\delta_{i-n-1}$ contains a $+1$ move or a $-1$ move for $\tau$, respectively.
@@
$\gamma^{0}_{i}=\begin{cases}\bar{\sigma}&\textrm{if}\;\bar{\phi}^{0}=\phi^{0}_{i}\\ b&\textrm{otherwise.}\end{cases}$
We now use a single attention layer to compute $\gamma^{\tau}_{i}$, the contents at $h^{\tau}_{i}$ on each non-input tape $\tau$. The layer uses two layer-norm hashes, again taking advantage of the multi-pre-norm architecture:
@@
$\psi^{\tau}_{i}$ $\triangleq\phi(1,1/i)=\phi(i,1).$
With strict causal masking, we attend with query $\langle\phi^{\tau}_{i},\vec{1}\rangle$, key $\langle\phi^{\tau}_{j},-\psi^{\tau}_{j}\rangle$, and value $\langle\phi^{\tau}_{j},\delta_{j-n-1}\rangle$. The head retrieves some (potentially averaged) value $\langle\bar{\phi}^{\tau},\bar{\delta}\rangle$ We will show that, if and only if there exists some $j<i$ such that $h^{\tau}_{j}=h^{\tau}_{i}$, this head retrieves one value from the greatest such $j$. In the forward direction, by Lemma 2, the first part of attention inner product is maximized at each $j$ such that $h^{\tau}_{i}=h^{\tau}_{j}$. The second component is maximized at the largest $j<i$ satisfying this condition. Thus, the head retrieves the value from the most recent update to cell $h^{\tau}_{i}$ of tape $\tau$ before step $i$, leveraging strict causal attention. We establish bidirectionality by proving the contrapositive. It must be that for all $j<i$, $h^{\tau}_{i}\succ h^{\tau}_{j}$, where $\succ$ is either $>$ or $<$. The head retrieves $\bar{\phi}^{\tau}=\frac{1}{\left\lvert M\right\rvert}\sum_{j\in M}\phi(h^{\tau}_{j},1)$ for some $M\subseteq\{1,\ldots,n\}$. Again, by Lemma 3, $\bar{\phi}^{\tau}\neq\phi^{\tau}_{h^{0}_{i}}$. Based on this property, we can compute the value on tape $\tau$ at $h^{\tau}_{i}$ as
$\gamma^{\tau}_{i}=\begin{cases}[\bar{\delta}]_{2+\tau}&\textrm{if}\;\bar{\phi}^{\tau}=\phi^{\tau}_{i}\\ b&\textrm{otherwise.}\end{cases}$
Having obtained all arguments of the transition function, we compute $\delta_{i-n}=\delta(q_{i-n-1},\sigma_{h^{0}_{i}},\gamma^{1}_{i},\ldots\gamma^{k+1}_{i})$ with a feedforward net.
@@
The critical role multi-pre-norm plays in Theorems 1 and 2 suggest it could be interesting to investigate incorporating multi-pre-norm into transformers in practice.
@@
This lets us reason about the power of transformer decoders in terms of Turing machines. Simulating an automaton (cf. Theorem 1) is not the only new capability unlocked with $\mathrm{O}(n)$ steps: rather, we can solve any problem a Turing machine can solve in $\mathrm{O}(n)$ time, such as simulating real-time counter machines (Weiss et al., 2018). With $\mathrm{O}(n^{2})$ steps, we can solve directed graph connectivity using standard graph traversal algorithms like depth-first search. Depth-first search runs in $\mathrm{O}(n)$ time on a random access Turing machine (Wigderson, 1992), which can be simulated in $\mathrm{O}(n^{2})$ time without random access. Possibly, transformers can solve directed graph connectivity with fewer than $\mathrm{O}(n^{2})$ steps, as results from Zhang et al. (2023) hint at.

### 11 Apr 2024 v5

@@ Abstract
Recent theoretical work has identified surprisingly simple reasoning problems, such as checking if two nodes in a graph are connected or simulating finite-state machines, that are provably unsolvable by standard transformers that answer immediately after reading their input. However, in practice, transformers’ reasoning can be improved by allowing them to use a “chain of thought” or “scratchpad”, i.e., generate and condition on a sequence of intermediate tokens before answering. Motivated by this, we ask: Does such intermediate generation fundamentally extend the computational power of a decoder-only transformer? We show that the answer is yes, but the amount of increase depends crucially on the amount of intermediate generation. For instance, we find that transformer decoders with a logarithmic number of decoding steps (w.r.t. the input length) push the limits of standard transformers only slightly, while a linear number of decoding steps, assuming projected pre-norm (a slight generalization of standard pre-norm), adds a clear new ability (under standard complexity conjectures): recognizing all regular languages. Our results also imply that linear steps keep transformer decoders within context-sensitive languages, and polynomial steps with generalized pre-norm make them recognize exactly the class of polynomial-time solvable problems—the first exact characterization of a type of transformers in terms of standard complexity classes. Together, this provides a nuanced framework for understanding how the length of a transformer’s chain of thought or scratchpad impacts its reasoning power.
@@ 1 Introduction
In this work, we address these open questions by characterizing the reasoning power of transformer decoders that can take intermediate steps before generating an answer and comparing them against transformers without intermediate steps. A transformer with a chain of thought constitutes a special case of a transformer decoder with intermediate steps. Our fine-grained results give upper and lower bounds on transformers’ power depending on $t(n)$: the number of allowed intermediate steps as a function of the input size $n$. We focus mainly on understanding three regimes: logarithmic steps (when $t(n)=\Theta(\log n)$), linear steps (when $t(n)=\Theta(n)$), and polynomial steps:
  1. 1.
@@
  3. 3.
    Linear Steps. Linear intermediate steps allow transformers with projected pre-norm333i.e., standard pre-norm but applied only to a linear projection of a sublayer’s input; cf. Definition 4 to simulate automata ($\mathsf{NC}^{1}$-complete), which cannot be done without intermediate steps unless $\mathsf{TC}^{0}=\mathsf{NC}^{1}$. Polynomial Steps. With a polynomial number of decoding steps, we show that transformers with strict causal attention and projected pre-norm are equivalent to the class $\mathsf{P}$. This, to our best knowledge, is the first equivalence between a class of transformers and a standard complexity class.
Together, our results provide a framework for understanding how the length of a transformer’s chain of thought affects its reasoning power. We find a logarithmic chain does not add much, while a linear chain affords more power on inherently sequential reasoning problems.
@@ 1.1 Main Results: Power of Transformers with Intermediate Decoding
Here $\mathsf{CoT}(t(n))$ denotes the set of languages recognized by some transformer using $t(n)$ decoding steps. Our lower bound (left side of Equation 1) assumes strict causal saturated attention and projected pre-norm, while upper bounds hold both with and without these architectural assumptions. Both our time lower bound and space upper bound are fairly tight: improving either by a factor larger than $\log t(n)$ would result in a fundamental complexity theory advance (Hopcroft et al., 1977).
@@
The left side of Equation 1 is proven by showing transformer decoders can simulate $t$ Turing machine steps with $t$ intermediate steps. Similar prior results have assumed a transformer with external memory (Schuurmans, 2023) or an encoder-decoder model with nonstandard-positional decodings (Pérez et al., 2021). Our construction adapts these ideas to work for a decoder-only model without external memory or extra positional encodings, but with strict causal masking and projected pre-norm (cf. Section 2.1). [^5] Our construction (Theorem 2) can be easily modified to work with an encoder-decoder model as well. The key idea behind our more general construction is the layer-norm hash (Section 3.1): a simple module for effectively storing memory in decoder-only transformers. We believe the layer-norm hash could be broadly useful for building algorithms in transformers. For example, Yao et al. (2021) used a related idea to construct transformers that recognize bounded-depth Dyck languages, although in a more ad hoc way.
@@ 2.1 Transformers
    For each $1\leq\ell\leq d$ and $1\leq k\leq h$, a head similarity function $s^{\ell}_{k}:\mathbb{D}_{p}^{m}\times\mathbb{D}_{p}^{m}\to\mathbb{D}_{p}$ whose form is defined in Section A.3 (and includes projected layer-norm);
@@
    For each $1\leq\ell\leq d$ and $1\leq k\leq h$, a head value function $v^{\ell}_{k}:\mathbb{D}_{p}^{m}\to\mathbb{D}_{p}^{m/h}$ whose form is defined in Section A.3 (and includes projected layer-norm);
@@
    For each $1\leq\ell\leq d$, an activation function $f^{\ell}:(\mathbb{D}_{p}^{m/h})^{h}\times\mathbb{D}_{p}^{m}\to\mathbb{D}_{p}^{m}$ whose form is defined in Section A.4 and implicitly uses the feedforward dimension $w$ (and includes projected layer-norm);
@@
    Projected Pre-Norm. Our lower bound constructions require $s_{\ell}$ and $f_{\ell}$ in Definition 2 to allow a generalization of standard pre-norm. Normally, a layer-norm is applied to the entire input to each sublayer. We generalize this, allowing each sublayer to apply a linear projection before layer-norm. Crucially, in particular, this enables each layer to pick out a subset of the previous hidden state to apply layer-norm to (cf. Definition 4 in Section A.1).
For convenience, our proofs with projected pre-norm use an even more general notion of pre-norm, namely multi-pre-norm, which allows each sublayer to take $k$ different projections of its input, apply layer-norm to each, and concatenate (cf. Definition 5 in Section A.1). Multi-pre-norm can, however, be simulated by multiple layers of projected pre-norm (see Section A.1 for a proof):
Proposition 1 (Chiang, 2024).
Multi-pre-norm with $k$ norms can be simulated by $k+1$ projected pre-norm layers.
@@ 2.2 Models of Computation
We define Turing machine computation in the standard way (cf. Appendix B).
@@ 3 Lower Bounds for Transformer Decoders
Prior work (Merrill & Sabharwal, 2023a) has established strong upper bounds on the reasoning problems transformers can solve. Specifically, under standard conjectures in complexity, transformers without intermediate decoding cannot recognize all regular languages. In this section, we show some of these shortcomings can be overcome with a suitable number of intermediate decoding steps (and projected pre-norm). Specifically, a linear number of steps enables simulating an automaton. We also show this can be extended to simulate $t(n)$ Turing machine steps with $t(n)$ decoding steps.
@@ 3.1 Introducing Layer-Norm Hash
$\phi(x,y)$ is a unit vector in $\mathbb{R}^{4}$. A key feature is scale invariance, and, in particular, that $\phi(x/i,1/i)$ is invariant w.r.t. $i$ in the sense that it is only a function of $x$, independent of $i$. Let $\phi_{x}\triangleq\phi(x,1)$. Then we have the following properties, whose proof may be found in Appendix C.
@@ 3.2 Simulating Automata
For any regular language $L$. there is a decoder-only projected pre-norm transformer with strict causal saturated attention (with or without positional encodings) that, on input \$$\$x$, [^7] Theorem 1 goes through without strict causal masking (but Theorem 2 will require strict ${}^{,}$ [^8] Theorems 1 and 2 both go through without \$currency-dollar\\$\$ as long as token $j$ can compute value $v_{j}=\mathbbm{1}[j=0]$. checks whether $x\in L$ with $\left\lvert x\right\rvert+1$ decoding steps.
@@
Let $A$ be a finite-state automaton recognizing $L$. We will simulate one step of $A$ with one transformer decoding step (after first reading $n$ input tokens). We refer to tokens with 00 indexing: \$currency-dollar\\$\$ is token 0, $x_{1}$ is token 1, etc. At step $i,n\leq i\leq 2n$, we will output a token $q_{i-n}$ encoding the next state of $A$. After printing the final state $q_{n}$, we use one additional step to output $1$ iff $q_{n}\in F$, the set of final states of $A$. At each token $i>0$, we compute $1/i$ by attending uniformly over the strict left context with value $\mathbbm{1}[x_{j}=\$]$. We show by induction that at step $i\geq n$, we can output $q_{i-n}$.
@@
Inductive Case: $i>n$. We already have a sequence of intermediate tokens $q_{0},\ldots,q_{i-n-1}$. Our goal is to compute $q_{i-n}=\delta(q_{i-n-1},\sigma_{i-n})$, which first involves retrieving $q_{i-n-1}$ and $\sigma_{i-n}$. $q_{i-n-1}$ is the input to the current column of the transformer. We will use hard attention to retrieve the current input symbol $\sigma_{i-n}$. To do this, we attend uniformly over the prior decoding tokens and \$, with a value of $1$ at \$ and 00 elsewhere. At tokens $i>n$ (i.e., decoding tokens), this yields $\frac{1}{i-n}$. Recall that projected pre-norm can simulate multi-pre-norm (Proposition 1). We now leverage the multi-pre-norm architecture to pass two layer-norms to a feedforward network:
@@
Theorem 1 shows that a linear number of decoding steps gives additional reasoning power to log-precision transformers with projected pre-norm (assuming $\mathsf{TC}^{0}\neq\mathsf{NC}^{1}$). This follows because log-precision transformers with no decoding steps are contained in uniform $\mathsf{TC}^{0}$ (Merrill & Sabharwal, 2023b), which means they cannot recognize all regular languages. In contrast, Theorem 1 says a linear number of steps is sufficient for recognizing all regular languages, establishing a conditional separation. This is an example of simple and familiar additional computational power granted by additional decoding steps. The core challenge in simulating an automaton is recurrence, which cannot be done without decoding steps (Merrill & Sabharwal, 2023b). A linear number of decoding steps allows simulating recurrence, which is where the additional power comes from. However, this added power does not stop with finite-state machines: the layer-norm hash can be used to simulate more complex models of computation such as Turing machines, which we will turn to next.
@@ 3.3 Simulating Turing Machines
We now show how a transformer decoder can simulate a Turing machine in real time using the layer-norm hash. Our decoder-only construction resembles the encoder-decoder construction of Pérez et al. (2021). However, it avoids simplifying assumptions from Pérez et al. (2021). In addition to assuming non-standard attention and no layer-norm, they required $1/i,1/i^{2},$ and $i$ in the positional embeddings, which is problematic because transformers cannot represent unbounded scalars like $i$ due to layer-norm. In contrast, our construction works with or without positional encodings. However, it assumes strict causal masking and projected pre-norm (Section 2.1).
@@
[…]
@@
Inductive Case: $i>n$. We first reconstruct $h^{\tau}_{i}$, the current position on each tape $\tau$. For each $\tau$, a head attends with query $1$, key $\mathbbm{1}[x_{j}\not\in\Sigma]$, and value being the move direction of $\tau$ at $j$. Since we assume strict causal attention (for reasons that will become clear later), head $\tau$ thus computes $h^{\tau}_{i-1}/i$. Since we need $h^{\tau}_{i}$, we write both $(h^{\tau}_{i-1}\pm 1)/i$ to the residual stream. When we need $h^{\tau}_{i}/i$ going forward, we use a linear layer to select either $(h^{\tau}_{i-1}+1)/i$ or $(h^{\tau}_{i-1}-1)/i$ depending on if the current input $\delta_{i-n-1}$ contains a $+1$ move or a $-1$ move for $\tau$, respectively.
@@
We now use a single attention layer to compute $\gamma^{\tau}_{i}$, the contents at $h^{\tau}_{i}$ on each non-input tape $\tau$. The layer uses two layer-norm hashes, again taking advantage of the multi-pre-norm architecture that projected pre-norm can simulate (Proposition 1):
@@
$\psi^{\tau}_{i}$ $\triangleq\phi(f(i),1),$
where $f(i)$ is defined in Definition 7 in Appendix E. Crucially, $f(i)$ is computable with a single transformer layer and monotonically decreasing with $i$. With strict causal masking, we attend with query $\langle\phi^{\tau}_{i},e_{1}\rangle$, key $\langle\phi^{\tau}_{j},-\psi^{\tau}_{j}\rangle$, and value $\langle\phi^{\tau}_{j},\delta_{j-n-1}\rangle$. Let $\langle\bar{\phi}^{\tau},\bar{\delta}\rangle$ be the head output. We show in Appendix E that two properties hold. First, by Lemma 6, $\bar{\phi}^{\tau}=\phi^{\tau}_{j}$ iff there is some $j<i$ s.t. $h^{\tau}_{i}=h^{\tau}_{j}$. Second, by Lemma 7, if there is some $j<i$ s.t. $h^{\tau}_{i}=h^{\tau}_{j}$, then the head retrieves $\langle\phi^{\tau}_{j},\delta_{j}\rangle$ for the greatest such $j$. Based on this, we compute the last-written value on tape $\tau$ at $h^{\tau}_{i}$ as $\gamma^{\tau}_{i}=[\bar{\delta}]_{2+\tau}$ if $\bar{\phi}^{\tau}=\phi^{\tau}_{i}$ and $\gamma^{\tau}_{i}=b$ otherwise. Having obtained all arguments for the transition function, we now compute $\delta_{i-n}=\delta(q_{i-n-1},\sigma_{h^{0}_{i}},\gamma^{1}_{i},\ldots\gamma^{k+1}_{i})$ with a feedforward net.
@@
The critical role projected pre-norm or multi-pre-norm play in Theorems 1 and 2 suggest it could be interesting to investigate incorporating these generalized pre-norms into transformers in practice.
Theorem 2 gives us a general result connecting the power of transformer decoders with $t(n)$ steps to Turing machines with the same number of intermediate steps:
@@
Thus, simulating an automaton (cf. Theorem 1) is not the only new capability unlocked with $\mathrm{O}(n)$ decoding steps: rather, such transformers can solve any problem a Turing machine can solve in $\mathrm{O}(n)$ time, such as simulating real-time counter machines (Weiss et al., 2018). With $\mathrm{O}(n^{2})$ steps, we can solve directed graph connectivity using standard graph traversal algorithms like depth-first search. Depth-first search runs in $\mathrm{O}(n)$ time on a random access Turing machine (Wigderson, 1992), which can be simulated in $\mathrm{O}(n^{2})$ time without random access. Possibly, transformers can solve directed graph connectivity with fewer than $\mathrm{O}(n^{2})$ steps, as results from Zhang et al. (2023) hint at.